OpenStudy (anonymous):

If x + |x| + y = 7 and x + |y| - y = 6 , then x + y =

OpenStudy (ksaimouli):

Okay, so is the answer 3?

OpenStudy (anonymous):

i need to know how to do it not the answer....

OpenStudy (anonymous):

When working with x + |x| + y = 7 and x = - 2 then y = 7 and when working withx + |y| - y = 6 and y = 2 then x = 6 So when x = -n then y = 7 and when y is not = - then x = 6 So with those restrictions we could say x + y = 7 + 6 = 13 Any thing else the solutions are limitless

OpenStudy (anonymous):

Where are the answers?

OpenStudy (anonymous):

I don't see them

OpenStudy (anonymous):

ksaimouli typed that it's 3 and yes it is but I need to know the method

OpenStudy (ksaimouli):

Here is my solution : for x>0 ; y>0 => x=6, y=-5 for x>0 ; y<0 => x=4, y=-1 => x+y = 3 (this is true because it fits the equation) for x<0 ; y>0 => x =6, y=7 for x<0, y<0 => y=7, x =20 Hence, x+y = 3

OpenStudy (anonymous):

Bit complicated

OpenStudy (ksaimouli):

@nixy that is also a solution, but doesn't fit into the equation

OpenStudy (anonymous):

but i got it